A capacitor consists of two closely spaced metal conductors of large area, separated by a thin insulating foil. It has an electrical apacity of 3800 V. Čalculate the amount of energy stored in the capacitor Submit Answer Tries 0/99 Calculate the charge on this capacitor when the electrical energy stored in the capacitor is 10.69 J Submit AwerTries 0/99 If the two plates of the capacitor have their separation increased by a factor of 5 while the charge on the plates remains constant, by 0 μF and is charged to a potential difference of 87.0 what factor is the energy stored in the capacitor increased?
Answer #1

1) Capacitance = 3800 uF

V = 87 V

for the energy stored

energy stored = 0.50 CV^2

energy stored = 0.50 * 3800 * 10^-6 * 87^2

energy stored = 14.38 J

2)

let the charge is Q

energy = Q^2/(2C)

10.69 = Q^2/(2 * 3800 *10^-6)

solving for Q

Q = 0.285 C

the charge on the capacitor is 0.285 C

3)

if the charge remains the same ,

when the seperation is increased , the capacitance will be 1/5 times

for the energy

energy = Q^2/(2C)

hence , energy will be 1/(1/5) times = 5 times

the energy on the capacitor will be 5 times